OpenStudy (anonymous):

What strategies do you use to convert a fraction to a decimal and a decimal into a fraction?

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

That is the question I was given.

OpenStudy (anonymous):

I don't have any examples. Sorry.:(

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

to turn a fraction in to a decimal you divide

OpenStudy (anonymous):

strategies

OpenStudy (anonymous):

?

OpenStudy (anonymous):

to turn a decimal in to a fraction a) if it is a terminating decimal, put it over the appropriate power of ten ex \[.12=\frac{12}{100}\]

OpenStudy (anonymous):

all i need is that strategies that are used just give me the names.

OpenStudy (anonymous):

to turn a fraction in to a decimal is not a "strategy" it is division \[\frac{3}{5}=3\div 5=0.6\] for example

OpenStudy (anonymous):

that is the question i got

OpenStudy (anonymous):

and that is the answer

OpenStudy (anonymous):

@ags2658 ???

OpenStudy (anonymous):

What is that strategy called?

OpenStudy (anonymous):

division

OpenStudy (anonymous):

which is actually a mathematical operation, but you can call it a "strategy" if you like to think of math as a war to be won

OpenStudy (anonymous):

trust me it is i am horrible @satellite73

OpenStudy (anonymous):

what about when you change it back @satellite73

OpenStudy (anonymous):

@pyxieyt

OpenStudy (anonymous):

i have an example above

OpenStudy (pyxieyt):

Yes?

OpenStudy (anonymous):

can you help

OpenStudy (pyxieyt):

@satellite73 Gave a great example, but ill try.

OpenStudy (anonymous):

All I need is the strats used

OpenStudy (pyxieyt):

To turn a fraction into a decimal you divide like satellite73 said above

OpenStudy (anonymous):

@sydneymmelin

OpenStudy (anonymous):

A fraction can be converted to its equivalent decimal by dividing its numerator by its denominator. For example, the fraction 3/4 can be converted to a decimal by dividing its numerator (3) by its denominator (4), which results in 0.75, its equivalent decimal. The fraction 5/8 can be similarly converted to a decimal by dividing 5 by 8, which results in 0.625.

OpenStudy (anonymous):

Can any of you help you are all on here.

OpenStudy (anonymous):

so is it division and multiplication

OpenStudy (anonymous):

division for converting a fraction to a decimal and multiplication for converting a decimal to a fraction

OpenStudy (anonymous):

Step 1: Write down the decimal divided by 1, like this: decimal/1 Step 2: Multiply both top and bottom by 10 for every number after the decimal point. (For example, if there are two numbers after the decimal point, then use 100, if there are three then use 1000, etc.) Step 3: Simplify (or reduce) the fraction

OpenStudy (anonymous):

you should copy this down so you remember.:)

OpenStudy (anonymous):

Sorry for taking so long, Aelia6th. I was busy with homework myself, and took a break. Well, when you say strategy, if mean ''trick'', there is one that applies to certain denominator. You must have a denominator of 10,100, 1000 and the like. If not, multiply the fraction by something to make sure you do. Take 5/100, for example. There are two zeros in the denominator. So, you take the numerator and move the decimal two places to the left. This creates .05, which is the decimal form of 5/100. Take 3/5 in your problem. You have a 5 in your denominator. In order for the ''shortcut''/strategy to work, you must change it to 10 (in other cases, it could have to changed to 100 or 1000). Simply multiply the numerator and denominator by 2: \[\frac{ 3 }{ 5 }(\times) \frac{ 2 }{ 2 } \] This gets us: \[\frac{ 6 }{ 10 }\] There is only one zero in the denominator. You take the numerator and move its decimal point one place to the left. This creates: \[0.6\] This is your answer. If this isn't what you were looking for, then I apologize. Besides this, I don't believe I remember anything else regarding fractions. I hope this helps! ;D